Tyrene Products manufactures recreational equipment. One of the companys products, a skateboard, sells for $37. The skateboa2. Due to an increase in labor rates, the company estimates that variable costs will increase by $1.85 per skateboard next ye5. Refer to the original data. The company is considering the construction of a new, automated plant. The new plant would sla0 Compute the degree of operating leverage. (Round your answer to 2 decimal places.) Degree of operating leverage
Answer #1

Solution 1:
a. Contribution margin ratio = Contribution margin / sales = $477,300 / $1591,000 = 30%
Contribution margin per unit = $37 - $25.90 = $11.10 per unit
Breakeven sales units = Fixed cost / contribution margin per unit = $277,500 / 11.10 = 25000 units
b. Degree of operating leverage = Contribution margin / Net operating income = $477,300 / $199,800 = 2.39


Solution 2:
New variable cost per unit =  $25.90 + $1.85 = $27.75 per ball
new contribution margin per unit = $37 - $27.75 = $9.25 per unit
New contribution margin ratio = $9.25 / $37 =25%
New breakeven point in balls = $277,500 / $9.25 = 30000 units


Solution 3:
Nos of balls to be sold to earn target income = (Fixed cost + Target profit) / contribution margin per unit
= ($277,500 + $199,800) / $9.25 = 51600 units

Solution 4:
Variable cost per unit = $27.75 per unit
Required contribution margin ratio = 30%
required variable cost ratio = 70%
New selling price per unit = $27.75 / 70% = $39.64 per unit


Solution 5:
New variable cost per unit = $25.90 * 80% = $20.72 per unit
New contribution margin per unit = 37- $20.72 = $16.28 per unit
New fixed costs = $277,500*192% = $532,800
New CM ratio = $16.28/$37 = 44%
New breakeven point = $532,800/ $16.28 = 32727 units

Solution 6a:
Nos of balls to be sold to earn target income = (Fixed cost + Target profit) / contribution margin per unit
= ($532,800 + $199,800) / $16.28 = 45000 units

Solution 6b:

Н 5 6 7 Tyrene Products 8 Degree of operating leverage Particulars Sales (43000 $37) Amount 9 1591000 10 Variable cost (43000